OpenStudy (anonymous):

15x+1= y solve for x

OpenStudy (anonymous):

\[x = \frac{y - 1}{15}\]

OpenStudy (anonymous):

can u show how to solve it

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[15x + 1 = y\]

OpenStudy (anonymous):

You want to try and get x on one side and all the other numbers on the other side. So the first thing im going to move is the 1. I want to get it to the other side. So to do that i would have to subtract 1 to get it to the other side like this.\[15x = y - 1\] Now we have 15x on one side and y - 1 on the other side. But we just want x on one side. So we need to get rid of the 15. Since 15 is being multiplied to x.. To move it to the other side we need to divide 15 on both sides and you end up with this \[x = \frac{y-1}{15}\]

OpenStudy (anonymous):

k